OpenStudy (anonymous):

Fans & Medals!! Solve the equation: -5(-4m – 8) = -2m - 26 Solve the equation: -5(-4m – 8) = -2m - 26 m = 11 m = -3 m = -11 m = 3

OpenStudy (anonymous):

distribute the -5 into everything in parenthesis (two negatives will make a positive) get M's on one side whole numbers on the other then divide to isolate m

OpenStudy (anonymous):

-5(-4m – 8) = -2m - 26 We need to distribute -5 into everything in the parenthesis. -5 * -4m = 20m -5 * -8 = 40 We're left with: 20m + 40 = -2m - 26 Add 2m to both sides to isolate m on one side. 20m + 40 = -2m - 26 + 2m +2m 22m + 40 = -26 Subtract 40 from both sides. 22m = -66 Lastly, divide 22 from each side. m = -3 <---- Answer. (:

OpenStudy (vanessaperezxo):

the answer is -3